OpenStudy (ihatemath):

Solve the equation. 8 – 2x = –8x + 14

OpenStudy (mattfort3212):

11/5

OpenStudy (fruitcakeninja):

Multiply first then add then take away

OpenStudy (lapinblanc):

8-2x=-8x+14 8(-2x+8x)=(-8x+8x)+14 8+6x=14 (8-8)+6x=(14+-8) 6x=6 6x/6=6/6 x=1

OpenStudy (fruitcakeninja):

PEMDAS

OpenStudy (yoloshroom):

@mattfort3212 Dont just give an answer babe Explain how you got it

OpenStudy (mattfort3212):

id rather just help out and if they ask how ill expalin hun. but ill try and do better.